A submarine obtains its fresh (pure) water supply by reverse osmosis. A sem-permeable membrane that allows only water to pass through is placed on the hull of the submarine. As the submarine descends into the ocean, the pressure inside is maintained at 1 atm while the pressure outside increases due to hydrostatic pressure. At a certain depth, this pressure difference across the membrane will become sufficient to cause water to diffuse into the submarine, against the water concentration gradient. The following information is available:
Answer #1

Mole fraction of the salt = 0.05

Basis : 100 mole of Ocean water

Moles of salt = 0.05*100 = 5 mol

Moles of water = (1-0.05)*100 = 95 mol

Mass of 95 mol of water = Moles * Molar mass = 95 mol * 18 g/mol = 1710 g

Volume of water = Mass of water / density = 1710 g/ 1 g/cm3 = 1710 cm3

Concentration of Salt in ocean = Moles of salt / Volume of water = 5/1710 mol/cm3 = 0.002924  mol/cm3 = 2924 mol/m3

Osmotic pressure (\Delta \pi) = \pi_D - \pi_F

\pi_D =  CRT = 2924 * 8.314 * T = 24310.136*T Pa

T = 290 - h/100 K

\pi_D = 24310.136* ( 290 - h/100) Pa

\pi_F = CRT = 0 . Assuming no salt or negligible amount of salt present in fresh water.

\Delta \pi = 24310.136* ( 290 - h/100) Pa

\Delta P = Pout - Pin

Pout = Patm + \rho _{sw} gh = 101325 + 1145 *9.81*h = 101325 + 11232.45*h Pa

Pin = 1 atm = 101325 Pa

\Delta P = Pout - Pin = 101325 + 11232.45*h - 101325 = 11232.45*h Pa

The water flux is given as,

Q = Lp*(\Delta P - \Delta \pi)

Just before water starts flowing in to low concentration area,

\Delta P - \Delta \pi = 0

On substituting the values of the above terms,

24310.136* ( 290 - h/100) = 11232.45*h

2.164*(290 - h/100) = h

h = 614.29 m

Minimum depth = 614.29 m